A typical family in an island economy consumes only apples, bananas and cloth. The base period is year 1. Prices in the base period are $2 per pound for applies. $3 per pound for bananas and $5 per yard of cloth. The typical family spends $40 on apples, $45 on bananas, $25 on cloth. In year 2, apples cost $3 per pound and cloth costs $8 per yard. Calculate the economy's Consumer Price Index for year 2 and the inflation rate between year 1 and year 2
Answer #1

Expenditure in Year 1
$40+45+25 = 110

Q of apples = 40/2 = 20

Q of bananas = 45/3=15

Q of cloth = 25/5=5

Expenditure in year 2 :

20x3+15x3+5x8=60+45+40 = 145

CPI in Year 2: 145/110x100 = 131.81

Inflation rate = Current CPI-Base year CPI/Base year CPI x 100

=131.81-100/100x100 = 31.8%